OpenStudy (anonymous):

Could someone tell me the steps of solving this 2nd order differential equation? (x^2)y'' - x y' + y = 8(x^3) Is it Euler's equations for the left side and method of undertermined coeffficients for the right side? Thanks!!!

OpenStudy (anonymous):

the A.E. is D^2 - D +1 =0 D= 1+- SQRT(3) i/ 2 yc= e^1/2 (C1 cos sqrt(3)/2 + C2 sin sqrt(3)/2)

OpenStudy (anonymous):

particular solution = 1/ (D^2 - D +1) * 8(x^3)

OpenStudy (anonymous):

sry an error in typing. yc= e^1/2 (C1 cos sqrt(3)/2 x + C2 sin sqrt(3)/2 x)

OpenStudy (anonymous):

ah again yc= e^1/2 x (C1 cos sqrt(3)/2 x + C2 sin sqrt(3)/2 x)

OpenStudy (anonymous):

why did you write: D^2 - D +1 =0 at the very beginning? Did you divide both sides by x squared?

OpenStudy (anonymous):

this is auxiliary equation

OpenStudy (anonymous):

yeah but you can't just ignore the x squared right? i know the quadratic formula but i am not sure if this is the case here

OpenStudy (anonymous):

still there buddy?

OpenStudy (anonymous):

ok give me a sec...

OpenStudy (anonymous):

so here we go...

OpenStudy (anonymous):

put z=logx and xD=theta , x^2D^2= theta (theta -1)

OpenStudy (anonymous):

1 sec

OpenStudy (anonymous):

Your way may work but I don't think I've seen it in class...it's gotta be easier than we think

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thanks a lot though

OpenStudy (anonymous):

you've been super helpful

OpenStudy (anonymous):

no prob... i wish i cud help u a lot more.. :(